. A quality controller wants to estimate the proportion of high quality goods out of a batch of products with a 90% confidence level and a margin of error of 5%. How many products must he test. 30


1
Expert's answer
2022-05-24T09:33:53-0400

A 90% confidence interval has a z-score (a critical value) "z_c=1.6449."


"ME=z_c\times\sqrt{\dfrac{p(1-p)}{n}}"

"=1.6449\times\sqrt{\dfrac{0.5(1-0.5)}{n}}=0.05"

"n=(16.449)^2"

"n=271"